57. Aleah is an electrical engineer. Her wage increased from $30 per hour to $40 per hour. She can work up to 50 hours each week. The table below shows her utility from different levels of leisure and income. If Aleah decreased her hours of work from 30 to 20 hours per week before her raise, the marginal utility loss from having less income is:

Hours of Leisure

Total Utility from Leisure

Income

Total Utility from Income

0

0

0

0

10

29

300

42

20

54

400

74

30

73

600

96

40

86

800

113

50

92

900

119

1200

122

1500

124

1600

126

2000

127

A. 23

B. 56

C. 19

D. 100

Answer: A Reference:
Answer #1

If she has decreased her hours of work from 30 to 20 hours per week, then her salary is reduced from 30*30 = 900 to 20*30 = 600. This represents a utility loss from 119 utils to 96 utils. Hence the marginal utility loss from having less income is 119 - 96 = 23 utils.

Select 23 utils. Option A.